# Multi-Knowledge-oriented Nighttime Haze Imaging Enhancer for Vision-driven Intelligent Transportation Systems

### v3

@@ V-C Real-world Degradation Analysis
TABLE IV: Ablation Studies on the TNL in ID, LLIE and NHIE tasks.
@@
TABLE V: Ablation Studies on the MRFE in ID, LLIE and NHIE tasks.
@@
TABLE VI: Ablation Studies on the Different Loss Function in ID, LLIE and NHIE tasks.
@@ V-D Ablation Study
V-D1 Ablation Studies on TNL
To validate the effectiveness of the proposed TNL modules, we assign different image restoration tasks to NHIE sub-node to train another version of MKoIE, as shown in Fig. 9, only a single sub-network was used to complete the three tasks. As shown in Table IV, compared to the full version of MKoIE, the performance of ID, LLIE, and NHIE degrades when the TNL module is specified independently. This highlights the effectiveness of the proposed shared and task-specific feature extraction methods. By leveraging its distinct architecture, the TNL module is designed to enhance feature extraction and thus improve overall task performance. It integrates the fundamental features required for ID, LLIE, and NHIE, without TNL module, MKoIE exhibits poor performance in the three tasks.
V-D2 Ablation Studies on Network Modules
To assess the significance of the MRFE module in MKoIE, we performed an ablation study by completely removing it. As shown in Table V, its removal consistently led to declines in PSNR and SSIM across all tasks. This demonstrates that the absence of the MRFE module limits the network’s ability to extract essential features, resulting in reduced restoration performance.
V-D3 Ablation Studies on Loss Function
Each component of the hybrid loss function show in subsection IV-F. To demonstrate the effectiveness of hybrid loss function, we train two versions: one using only the reconstruction loss and the other using only the perceptual loss. From Table VI, the declines in both PSNR and SSIM indicate that hybrid loss function guide the network toward improved restoration performance. Reconstruction and perceptual loss enhances restoration performance by balancing numerical stability, outlier resilience, perceptual consistency, and semantic preservation, achieving optimal results.
V-E Improving SOD Tasks with MKoIE
TABLE VII: The comparison for SOD task between original images and images enhanced by the MKoIE includes MAE, max F-measure($F_{\beta}^{\text{max}}$), max E-measure($E_{\phi}^{\text{max}}$)and S-measure($S_{\alpha}$).
@@ V-E Improving SOD Tasks with MKoIE
We utilized U2Net [48] and show its performance on CDD dataset. Notably, we directly applied the pretrained U2Net model without further training, achieving a marked enhancement in SOD through MKoIE preprocessing. As shown in Fig. 10 and Table VII, the U2Net model achieved significant improvements in MKoIE-processed images, enabling more accurate detection and segmentation of salient objects. This result validates the effectiveness of MKoIE, demonstrating its practical application value in traffic scenarios under complex weather conditions. Specifically, in autonomous driving and traffic monitoring systems, it enhances the detection and segmentation performance of critical targets such as pedestrians and vehicles, thereby further improving traffic safety.
V-F Complexity and Running Time Comparison
TABLE VIII: We selected the top four models based on the quantitative comparison of the NHIE task and present the corresponding average runtime (in seconds).
@@ V-F Complexity and Running Time Comparison
MKoIE adopts an efficient structural design, which can significantly improve the image recovery speed in low light and haze environments. We randomly selected 50 images from the test set and conducted tests under a consistent hardware setup. As shown in Table VIII, the average processing time of our model is about 1.68 seconds. The results show that MKoIE has advantages in both efficiency and effectiveness, demonstrating its potential for deployment in time-sensitive ITS scenarios such as traffic monitoring and autonomous driving.
VI Conclusions
This paper proposes a multi-knowledge-oriented imaging enhancer (MKoIE) to address salient object detection challenges in ITS under adverse imaging conditions. MKoIE effectively handles three types of degradation (i.e., daytime haze, low light, and nighttime haze) through a task-oriented node learning mechanism and self-attention module, while achieving efficient multi-scale feature extraction through a multi-receptive field enhancement module, making the method highly suitable for real-time ITS applications. Experimental results demonstrate that the proposed MKoIE achieves superior performance across various weather and imaging conditions, thus providing an effective solution for image enhancement and salient object detection in challenging environments, contributing to safer and more reliable transportation systems.

### v4

@@ 5.3 Real-world Degradation Analysis
Table 4: Ablation Studies on the TNL in ID, LLIE and NHIE tasks.
@@
Table 5: Ablation Studies on the MRFE in ID, LLIE and NHIE tasks.
@@
Table 6: Ablation Studies on the Different Loss Function in ID, LLIE and NHIE tasks.
@@ 5.4 Ablation Study
5.4.1 Ablation Studies on TNL
To validate the effectiveness of the proposed TNL modules, we assign different image restoration tasks to NHIE sub-node to train another version of MKoIE, as shown in Fig. 9, only a single sub-network was used to complete the three tasks. As shown in Table 4, compared to the full version of MKoIE, the performance of ID, LLIE, and NHIE degrades when the TNL module is specified independently. This highlights the effectiveness of the proposed shared and task-specific feature extraction methods. By leveraging its distinct architecture, the TNL module is designed to enhance feature extraction and thus improve overall task performance. It integrates the fundamental features required for ID, LLIE, and NHIE, without TNL module, MKoIE exhibits poor performance in the three tasks.
5.4.2 Ablation Studies on Network Modules
To assess the significance of the MRFE module in MKoIE, we performed an ablation study by completely removing it. As shown in Table 5, its removal consistently led to declines in PSNR and SSIM across all tasks. This demonstrates that the absence of the MRFE module limits the network’s ability to extract essential features, resulting in reduced restoration performance.
5.4.3 Ablation Studies on Loss Function
Each component of the hybrid loss function show in subsection 4.6. To demonstrate the effectiveness of hybrid loss function, we train two versions: one using only the reconstruction loss and the other using only the perceptual loss. From Table 6, the declines in both PSNR and SSIM indicate that hybrid loss function guide the network toward improved restoration performance. Reconstruction and perceptual loss enhances restoration performance by balancing numerical stability, outlier resilience, perceptual consistency, and semantic preservation, achieving optimal results.
5.5 Improving SOD Tasks with MKoIE
Table 7: The comparison for SOD task between original images and images enhanced by the MKoIE includes MAE, max F-measure($F_{\beta}^{\text{max}}$), max E-measure($E_{\phi}^{\text{max}}$)and S-measure($S_{\alpha}$).
@@ 5.5 Improving SOD Tasks with MKoIE
We utilized U2Net [52] and show its performance on CDD dataset. Notably, we directly applied the pretrained U2Net model without further training, achieving a marked enhancement in SOD through MKoIE preprocessing. As shown in Fig. 10 and Table 7, the U2Net model achieved significant improvements in MKoIE-processed images, enabling more accurate detection and segmentation of salient objects. This result validates the effectiveness of MKoIE, demonstrating its practical application value in traffic scenarios under complex weather conditions. Specifically, it enhances the detection and segmentation of critical visual targets, thereby improving environmental perception and supporting reliable decision-making in autonomous visual systems.
5.6 Complexity and Running Time Comparison
Table 8: We selected the top four models based on the quantitative comparison of the NHIE task and present the corresponding average runtime (in seconds).
@@ 5.6 Complexity and Running Time Comparison
MKoIE adopts an efficient structural design, which can significantly improve the image recovery speed in low light and haze environments. We randomly selected 50 images from the test set and conducted tests under a consistent hardware setup. As shown in Table 8, the average processing time of our model is about 1.68 seconds. The results show that MKoIE demonstrates advantages in both efficiency and effectiveness, demonstrating its potential for deployment in time-sensitive and real-world intelligent imaging applications.
6 Conclusions
This paper proposes a multi-knowledge-oriented imaging enhancer (MKoIE) to address salient object detection challenges in VIS under adverse imaging conditions. MKoIE effectively handles three types of degradation (i.e., daytime haze, low light, and nighttime haze) through a task-oriented node learning mechanism and self-attention module, while achieving efficient multi-scale feature extraction through a multi-receptive field enhancement module, making the method highly suitable for real-time VIS applications. Experimental results demonstrate that the proposed MKoIE achieves superior performance across various weather and imaging conditions, thus providing an effective solution for image enhancement and salient object detection in challenging environments, contributing to safer and more reliable intelligent systems.
Acknowledgements
The work described in this paper is supported by the Research Grants Council of the Hong Kong Special Administrative Region, China (Grant Number. PolyU 15201722), the Key Research and Development Project of Sichuan (Grant Number. 2024YFG0008), and the Major Program of National Natural Science Foundation of China (Grant Number. T2293771).